OpenStudy (anonymous):

What is the interquartile range (IQR) for the data set? 12, 14, 15, 18, 22, 24, 28, 31, 35

OpenStudy (anonymous):

IQR = Q3 - Q1

OpenStudy (anonymous):

Q3 = 28 Q1 = 15 Can you figure it out?

OpenStudy (anonymous):

@ashhoperawr

OpenStudy (anonymous):

im triying to im not very good at the IQR everything else im fine with its jhust a lil confusing from what everything is saying

OpenStudy (anonymous):

would the answer be 23?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

Our equation is IQR = Q3 - Q1

OpenStudy (anonymous):

Q3 is 28 and Q1 is 15. So We have IQR = 28 - 15

OpenStudy (anonymous):

13

OpenStudy (anonymous):

a. 14.5 b. 15 c.23 d. 29.5

OpenStudy (anonymous):

Ah, hold on

OpenStudy (anonymous):

The answer is B, 15.

OpenStudy (anonymous):

Q3 is 29.5, and Q1 is 14.5

OpenStudy (anonymous):

so its not A